OpenStudy (anonymous):

Find dy/dx of (1+cos^2 7x)^3

OpenStudy (anonymous):

chain rule for this one

OpenStudy (anonymous):

\[\frac{d}{dx}(1+\cos^2(7x))^3=3(1+\cos^2(7x))^2\times (2\cos(7x))\times (-\sin(7x))\times 7\]clean up with algebra

OpenStudy (anonymous):

So, How do you find d/dx of cos^2(x)? I can't quite remember if it is just 2cos(x), or if you have to do something special to it.

OpenStudy (anonymous):

chain rule first take care of the square, then take care of the cosine

OpenStudy (anonymous):

the derivative of \(f^2(x)\) is \(2f(x)f'(x)\)

OpenStudy (anonymous):

So, it will become 2cos(x), then become -2sin(x)?

OpenStudy (anonymous):

no the derivative of some function squared is two time that function, times the derivative of that function more succinctly written as \[(f^2)'=2ff'\]

OpenStudy (anonymous):

So you use the product rule?